OpenStudy (anonymous):

find the equation of the line tangent to y=arctanx at x=1

OpenStudy (anonymous):

y=arctanx at x=1 y '= 1/(1+x^2) y '(1)=1/2=0.5 at x=1,y=pi/4 y-y1=m(x-x1) y-pi/4=0.5x-0.5 y=0.5x-0.5+pi/4

OpenStudy (anonymous):

A solution plot is attached. To eliminate the fractions, the line is labeled in Standard Form.